# the Philippine Eagle, and cosmological parameter estimation

### v2

@@ 4.4 Philippine Eagle Optimization Algorithm
Begin Input objective function $f$, dimension $D$, search space bounds $X_{\textrm{min}}$, $X_{\textrm{max}}$ Initialize maximum function evaluations $N_{\textrm{max}}$, initial eagle population size $S_{0}$, local food size $S_{\textrm{loc}}$ Set equal probabilities of the three operators Movement, Mutation I, and Mutation II and initialize memory elements for the scaling factors $F$ Generate initial population of eagles $X$ and select the best eagle $X^{\star}$ of the population based on function value Conduct local search for the best food $Y^{\star}$ of the best eagle using interior point method Update eagle population size $S$ using linear population size reduction Divide eagle population into three subpopulations using the probabilities of the three operators Apply Movement operator to the first subpopulation of eagles: $(X_{\textrm{new}})_{i}=X_{i}+F_{i}\cdot(X^{\star}-X_{i}\,+\,X_{r_{1}}-X_{\textrm{arc}}\,+e^{-d^{2}}\cdot(X_{\textrm{near}}-X_{i}))$ Apply Mutation I operator to the second subpopulation of eagles: $(X_{\textrm{new}})_{i}=F_{i}\cdot(X_{r_{1}}+{X^{\star}-X_{r_{2}})}+S\cdot L(D)$ Apply Mutation II operator to the third subpopulation of eagles: $(X_{\textrm{new}})_{i}=F_{i}\cdot(\hat{X}+X^{\star}-X_{\textrm{mean}})$ Generate new population of eagles by replacing old eagles with the new eagles $X_{\textrm{new}}$ that gained improvement in function value Update the best eagle $X^{\star}$ of the population and search for the new best food $Y^{\star}$ of the updated best eagle using interior point method Is the function value error $|f(Y^{\star})-f_{\textrm{true}}|$ less than $10^{-8}$? Is the current number of function evaluations $N$ greater that the maximum function evaluations? Output searched food $Y^{\star}$ of the best eagle and its function value $f(Y^{\star})$ End Update the probabilities of the three operators Movement, Mutation I, and Mutation II based on the improvement rate of each operator Update the memory elements for the scaling factors using the weighted Lehmer mean of the scaling factors of improved eagles YesNoYesNo
@@
As with GA and IMODE, we take the PEOA objective function to be the likelihoods (4.2-4.4) and generate the uncertainty by applying the method iteratively to bootstrapped data. We use default configurations with PEOA, with a minimum eagle population size of 5, cluster size factor 0.04, and Max-FES 500 [47]. Our detailed implementation of PEOA can be found in github.com/ErikaAntonette/Philippine-Eagle-Optimization-Algorithm.
@@ 4.5 Uncertainty Analysis
0:  Mock data $(z_{\rm mock},y_{\rm mock})$, standard deviation $\Delta y_{\rm mock}$, number of realizations $N$
@@
7:     Optimize fitness functions in Eqs. (4.24.4) using a metaheuristic algorithm
@@
Incorporating uncertainty quantification into the optimization framework of a metaheuristic algorithm allows for analysis of the statistical behavior of the estimated parameters. Given the mock data $(z_{\rm mock},y_{\rm mock})$, we introduce normally distributed noise to $y_{\rm mock}$ with $\Delta y_{\rm mock}$ as the standard deviation, effectively generating perturbed realizations of the observed data. This process is repeated ${\cal O}\left(10^{3-4}\right)$ times, creating a comprehensive set of data with uncertainty. For each of these data sets, we perform optimization using GA, IMODE, and PEOA on the fitness functions described in Eqs. (4.24.4). The resulting optimizers from each iteration are stored and analyzed to characterize their statistical distributions, providing assessment into the reliability and variability of the obtained parameters under uncertainty. This is summarized in Algorithm 1.
@@ 5.1 Expansion Rate + Growth Rate
We examine the posterior estimates obtained when considering only expansion and growth rate observations. For each of the 500 mock data sets analyzed, posteriors were derived using MCMC, GA, IMODE, and PEOA.
@@
${\rm RDEV}[a]=\dfrac{\overline{a}-a^{\rm true}}{a^{\rm true}}\,,$ (5.1)
where $\overline{a}$ represents the sample mean and $a^{\rm true}$ the true input parameters, i.e., $\Omega_{m0}=0.3$, $H_{0}=70$ km s-1Mpc-1, and $\sigma_{8}=0.8$. Precision is defined as:
${\rm PREC}[a]=\dfrac{{\rm STD}[a]}{{\rm AVE}[a]}=\dfrac{\Delta a}{\overline{a}}\,,$ (5.2)
measuring normalized uncertainty. Figure 7 shows the accuracy and precision for 500 realizations.
@@
Precision, on the other hand, differs across methods, as shown in the bottom panel of Figure 7. GA is the least precise, followed by IMODE, while PEOA and MCMC demonstrate higher precision. This distinction aligns with Figure 6, where the precision order can be visually inferred from the contour sizes. For $\Omega_{m0}$, average precision values are approximately 11%, 23%, 17%, and 11% for MCMC, GA, IMODE, and PEOA, respectively. Similar trends are observed for $H_{0}$ and $\sigma_{8}$, highlighting consistent precision hierarchies across parameters and covariances.
@@
$D_{0}(f(Z)|g(Z))=\dfrac{1}{N_{Z}}\sum_{Z}\left(\dfrac{f(Z)-g(Z)}{\sqrt{\Delta f(Z)^{2}+\Delta g(Z)^{2}}}\right)^{2}\,,$ (5.3)
@@
Figure 8: Histograms for the goodness (5.3) of the reconstructed cosmological functions: the expansion rate and the growth rate.
@@ 5.2 Expansion Rate + Growth Rate + Supernovae
Figure 11: Histograms for the goodness (5.3) of the reconstructed expansion rate, growth rate, and distance modulus.
Despite the differences highlighted by accuracy and precision metrics, the goodness-of-reconstruction metric (5.3) (Figure 11) shows comparable results for all methods. This metric, while valuable, does not fully capture the discrepancies between methods revealed by other metrics. IMODE and GA exhibit tendencies to overfit the expansion and growth rate data relative to MCMC and PEOA, which produce nearly indistinguishable histograms. All methods appear to overfit the SNe data, indicating the need to assess reconstructions using diverse evaluation metrics [20, 75].
@@ 6 Discussion
Table 1 summarizes the ensemble averages (over 500 mock data sets) of the accuracy and precision of MCMC, GA, IMODE, and PEOA in $\Lambda$CDM cosmological parameter estimation. Note that these statistics are the same as Figures 7 and 10.
Table 1: Ensemble averages (multiple realizations) of accuracy (Eq. (5.1)) and precision (Eq. (5.2)) of MCMC, GA, IMODE, and PEOA for $\Lambda$CDM parameter estimation. $H(z)$+$f\sigma_{8}(z)$ refers to when only expansion rate and growth rate measurements are considered; and $H(z)$+$f\sigma_{8}(z)$+$\mu(z)$ when supernovae is additionally considered.
Data set Method RDEV[$\Omega_{m0}$] PREC[$\Omega_{m0}$] RDEV[$H_{0}$] PREC [$H_{0}$] RDEV[$\sigma_{8}$] PREC[$\sigma_{8}$]
$H(z)$+$f\sigma_{8}(z)$ MCMC $-17\%$ $11\%$ $3\%$ $3\%$ $5\%$ $3\%$
GA $-11\%$ $23\%$ $2\%$ $6\%$ $3\%$ $7\%$
IMODE $-14\%$ $17\%$ $3\%$ $5\%$ $4\%$ $5\%$
PEOA $-17\%$ $11\%$ $3\%$ $3\%$ $5\%$ $3\%$
$H(z)$+$f\sigma_{8}(z)$+$\mu(z)$ MCMC $-5\%$ $5\%$ $0.2\%$ $0.3\%$ $0.7\%$ $2\%$
GA $5\%$ $53\%$ $-2\%$ $4\%$ $-6\%$ $15\%$
IMODE $16\%$ $39\%$ $-0.5\%$ $2\%$ $-4\%$ $11\%$
PEOA $-13\%$ $9\%$ $0.8\%$ $0.5\%$ $3\%$ $3\%$
With expansion and growth rate data ($H(z)$+$f\sigma_{8}(z)$), all methods performed comparably well. However, when supernovae data ($\mu(z)$) was added, PEOA consistently outperformed IMODE and GA, matching the performance of MCMC.
@@
To further investigate the performance of EAs, we analyzed their behavior for a single mock data realization of expansion rate, growth rate, and supernovae measurements, varying the hyperparameter tied to the maximum number of function evaluations (Max-FES). Max-FES, determined by the product of the population size and the number of generations, limits the total number of objective function evaluations an EA can perform (Section 4). While a standard Max-FES value of 500 sufficed for accurate and precise parameter estimation with $H(z)$+$f\sigma_{8}(z)$ data, this setting proved inadequate for IMODE and GA with default settings when SNe data were added. Table 2 presents the precision obtained by each method and the corresponding computational time for varying Max-FES values.
@@
Method Max-FES PREC[$\Omega_{m0}$] PREC [$H_{0}$] PREC[$\sigma_{8}$] Time (s)
@@
It is worth noting that the results of each method are consistent in all cases, though the constraints are not centered relative to the expected values [^6] This can be addressed by including the covariance in the analysis, since our mock data sets are based on correlated observations that were treated independently.. Regardless, the overall offset is consistent for all of the methods and for this work important observation is the performance of EA with respect to MCMC. The whisker plot supports the sentiments on accuracy and precision of EA that have been reiterated throughout the results and the discussion.
@@ 7 Conclusions
In this work, we introduced two novel evolutionary algorithms (EAs) for cosmological parameter estimation: the Improved Multi-Operator Differential Evolution (IMODE) and the Philippine Eagle Optimization Algorithm (PEOA). Among these, PEOA demonstrated performance comparable to the standard Markov Chain Monte Carlo (MCMC) method while surpassing other EAs, such as IMODE and the genetic algorithm (GA). These results underscore the potential of PEOA in cosmological applications, although it is crucial to interpret such findings within the specific context of parameter estimation. In optimization problems, there is no universally superior method; thus, this study serves as an initial benchmark for exploring EAs in cosmology.
@@
The motivation for exploring alternative statistical methods is not to replace existing techniques but to complement them. This is particularly relevant in cosmology, where diverse data sets—ranging from supernovae and the cosmic microwave background to baryon acoustic oscillations and gravitational waves—must be reconciled under a consistent physical framework. Similarly, the application of diverse statistical methods, including EAs, can provide robust cross-validation and novel perspectives. Our work represents a step toward establishing a more diverse set of tools for confronting cosmological data.
Future research directions naturally emerge from this study. First, while GA has been applied to real cosmological data, the potential of IMODE and PEOA, among a full universe of nature-inspired EAs, remains to be tested on such data sets, which include full covariance structures. This will offer insights into their performance in more realistic scenarios. Additionally, EAs could play a significant role in cosmological model selection and in reconstructing cosmological functions such as spatial curvature and the dark energy density. Another promising avenue lies in leveraging the ability of EAs to explore high-dimensional and complex parameter spaces to investigate potential systematic effects in the data. On a practical side, we are preparing user-friendly python implementations of IMODE and PEOA to facilitate their adoption to cosmology and astrophysics.
Acknowledgments
@@ Appendix A Pseudo-algorithms
  • $[X_{\textrm{min}},X_{\textrm{max}}]$: search space
@@ A.1 Genetic Algorithm
  • $PS$: population size
  • $cr$: crossover rate
@@
0:  $f$, $D$, $PS$, lower bound $\mathbf{X}_{\min}$, upper bound $\mathbf{X}_{\max}$, $cr$, $m$
0:  $f(\mathbf{x}^{*})$, $\mathbf{x}^{*}$
1:  Generate initial random population of size $PS$.;
@@
17:     The new generation is the collection of the elites and the offspring while maintaining the population size $PS$.
@@ A.2 Improved Multi-Operator Differential Evolution Algorithm
1:  Define $Prob_{ls}\leftarrow 0.1$, $Max\text{-}FES$, $prob_{1}\leftarrow 1$, $prob_{2}\leftarrow 1$, $NP$, $G\leftarrow 1$, and $FES\leftarrow 0$;
2:  Generate an initial random population $(X)$ of size $NP$;
@@
7:     Generate new population using the assigned DE operators, i.e., each operator $op$ evolves its assigned number of individuals $NP_{op}$;
8:     Calculate the diversity obtained from each operator $op$ $\left(D_{op}\right)$ and the quality rate of solutions $QR_{op}$;

### v3

@@ 4.4 Philippine Eagle Optimization Algorithm
Begin Input objective function $f$, dimension $D$, search space bounds $X_{\textrm{min}}$, $X_{\textrm{max}}$ Initialize maximum function evaluations $N_{\textrm{max}}$, initial eagle population size $S_{0}$, local food size $S_{\textrm{loc}}$ Set equal probabilities of the three operators Movement, Mutation I, and Mutation II and initialize memory elements for the scaling factors $F$ Generate initial population of eagles $X$ and select the best eagle $X^{\star}$ of the population based on function value Conduct local search for the best food $Y^{\star}$ of the best eagle using interior point method Update eagle population size $S$ using linear population size reduction Divide eagle population into three subpopulations using the probabilities of the three operators Apply Movement operator to the first subpopulation of eagles: $(X_{\textrm{new}})_{i}=X_{i}+F_{i}\cdot(X^{\star}-X_{i}\,+\,X_{r_{1}}-X_{\textrm{arc}}\,+e^{-d^{2}}\cdot(X_{\textrm{near}}-X_{i}))$ Apply Mutation I operator to the second subpopulation of eagles: $(X_{\textrm{new}})_{i}=F_{i}\cdot(X_{r_{1}}+{X^{\star}-X_{r_{2}})}+S\cdot L(D)$ Apply Mutation II operator to the third subpopulation of eagles: $(X_{\textrm{new}})_{i}=F_{i}\cdot(\hat{X}+X^{\star}-X_{\textrm{mean}})$ Generate new population of eagles by replacing old eagles with the new eagles $X_{\textrm{new}}$ that gained improvement in function value Update the best eagle $X^{\star}$ of the population and search for the new best food $Y^{\star}$ of the updated best eagle using interior point method Is the function value error $|f(Y^{\star})-f_{\textrm{true}}|$ less than $10^{-8}$? Is the current number of function evaluations NN greater that the maximum function evaluations? Output searched food $Y^{\star}$ of the best eagle and its function value $f(Y^{\star})$ End Update the probabilities of the three operators Movement, Mutation I, and Mutation II based on the improvement rate of each operator Update the memory elements for the scaling factors using the weighted Lehmer mean of the scaling factors of improved eagles YesNoYesNo
@@
Because PEOA is inspired by the Philippine Eagle, several terms in the algorithm are derived from this analogy. For example, an “eagle” is a candidate solution to the given optimization problem, similar to a “chromosome” in GA. The eagle with the least function value (in a minimization problem) is called the “best eagle” of the eagle population. The best eagle then conducts a local search, which is the use of an optimizer (e.g., the interior point method) to find the minimum point in a small region containing the best solution. In PEOA, the minimum point found by the best eagle during its local search is its “food,” and consequently, the local search is also called the “local food search.” If the stopping criterion has not yet been satisfied, PEOA proceeds to generate a new eagle population using three operators, namely “Movement,” “Mutation I,” and “Mutation II” (see Figure 5). These operators work in a similar manner as the crossover, mutation, and selection operators of GA.
As with GA and IMODE, we take the PEOA objective function to be the likelihoods (8-10) and generate the uncertainty by applying the method iteratively to bootstrapped data. We use default configurations with PEOA, with a minimum eagle population size of 5, cluster size factor 0.04, and Max-FES 500 [48]. Our detailed implementation of PEOA can be found in github.com/ErikaAntonette/Philippine-Eagle-Optimization-Algorithm. Lastly, we use the default parameter values of the PEOA adopted in its main reference paper [48], which are listed in Table 4.4.
Table 4: The default values of the hyperparameters of PEOA that are used in the simulations.
Parameter Value
dimension 3
maximum number of function evaluations 500
initial population size 180
minimum population size 5
maximum number of function evaluations during local search 90
cluster size factor 0.04
archive rate 2.6
memory size 60
@@ 4.5 Uncertainty Analysis
0: Mock data $(z_{\rm mock},y_{\rm mock})$, standard deviation $\Delta y_{\rm mock}$, number of realizations NN
@@
7:  Optimize fitness functions in Eqs. (810) using a metaheuristic algorithm
@@
Incorporating uncertainty quantification into the optimization framework of a metaheuristic algorithm allows for analysis of the statistical behavior of the estimated parameters. Given the mock data $(z_{\rm mock},y_{\rm mock})$, we introduce normally distributed noise to $y_{\rm mock}$ with $\Delta y_{\rm mock}$ as the standard deviation, effectively generating perturbed realizations of the observed data. This process is repeated ${\cal O}\left(10^{3-4}\right)$ times, creating a comprehensive set of data with uncertainty. For each of these data sets, we perform optimization using GA, IMODE, and PEOA on the fitness functions described in Eqs. (810). The resulting optimizers from each iteration are stored and analyzed to characterize their statistical distributions, providing assessment into the reliability and variability of the obtained parameters under uncertainty. This is summarized in Algorithm 1.
We emphasize that the uncertainty obtained using the above described bootstrap mechanism is conceptually different compared to ones estimated by a Bayesian method like MCMC. The bootstrap approximates frequentist confidence intervals, which quantify parameter variations under repeated sampling of the data. In contrast, Bayesian intervals represent regions of parameter space with a fixed posterior probability, given the observed data and prior information.
@@ 5.1 Expansion Rate + Growth Rate
We examine the posterior estimates obtained when considering only expansion and growth rate observations. For each of the 500 mock data sets analyzed, posteriors were derived using MCMC, GA, IMODE, and PEOA. The results were computed using a single machine to guarantee computational budget fairness across methods.
@@
${\rm RDEV}[a]=\dfrac{\overline{a}-a^{\rm true}}{a^{\rm true}}\,$, (13)
where $\overline{a}$ represents the sample mean and $a^{\rm true}$ the true input parameters, i.e., $\Omega_{m0}=0.3$, $H_{0}=70$ km s-1Mpc-1, and $\sigma_{8}=0.8$. When RDEV = 0, then the sample mean lies exactly on the true values. When RDEV is positive (negative), the sample mean lies to the right (left) of the true values. Precision is defined as:
${\rm PREC}[a]=\dfrac{{\rm STD}[a]}{{\rm AVE}[a]}=\dfrac{\Delta a}{\overline{a}}\,$, (14)
measuring normalized uncertainty. This purely positive quantity measures the width of a sample relative to its mean value. In general, small RDEV and PREC correspond to accurate and precise estimates, respectively. The question of how small or large depends on the sample and the objectives of the work. In our case we find that these two metrics are sufficiently able to distinguish the methods under study. Of course, alternative metrics are available, but we shall leave this for future studies. Figure 7 shows the accuracy and precision for 500 realizations.
@@
Precision, on the other hand, differs across methods, as shown in the bottom panel of Figure 7. GA is the least precise, followed by IMODE, while PEOA and MCMC demonstrate higher precision. This distinction aligns with Figure 6, where the precision order can be visually inferred from the contour sizes. For $\Omega_{m0}$, average precision values are approximately 11%, 23%, 17%, and 11% for MCMC, GA, IMODE, and PEOA, respectively. Similar trends are observed for $H_{0}$ and $\sigma_{8}$, highlighting consistent precision hierarchies across parameters and covariances. Note that the bottom panel of Figure 7 (PREC histograms) additionally attests to the stability of the bootstrap interval across realizations; in the sense that the histograms of the EAs can be visualized within compact distribution functions with well defined moments.
@@
$D_{0}(f(Z)|g(Z))=\dfrac{1}{N_{Z}}\sum_{Z}\left(\dfrac{f(Z)-g(Z)}{\sqrt{\Delta f(Z)^{2}+\Delta g(Z)^{2}}}\right)^{2}\,$, (15)
@@
Figure 8: Histograms for the goodness (15) of the reconstructed cosmological functions: the expansion rate and the growth rate.
@@ 5.2 Expansion Rate + Growth Rate + Supernovae
Figure 11: Histograms for the goodness (15) of the reconstructed expansion rate, growth rate, and distance modulus.
Despite the differences highlighted by accuracy and precision metrics, the goodness-of-reconstruction metric (15) (Figure 11) shows comparable results for all methods. This metric, while valuable, does not fully capture the discrepancies between methods revealed by other metrics. IMODE and GA exhibit tendencies to overfit the expansion and growth rate data relative to MCMC and PEOA, which produce nearly indistinguishable histograms. All methods appear to overfit the SNe data, indicating the need to assess reconstructions using diverse evaluation metrics [20, 83].
@@ 6 Discussion
[…]
Data set Method RDEV[$\Omega_{m0}$] PREC[$\Omega_{m0}$] RDEV[$H_{0}$] PREC $H_{0}$] RDEV[$\sigma_{8}$] PREC[$\sigma_{8}$]
[…]
GA $11\%$ $3\% 22\%$ $6\%$ $3\%$ $7\%$
IMODE $14\%$ $17\%$ $3\%$ $5\%$ $4\%$ $5\%$
PEOA $7\% 11$$1\%$ $\% 33\%$ $5\%$ $3\%$
[…]
GA $5\%$ $53\%$ $2\%$ $4\%$ $6\%$ $15\%$
IMODE $16\%$ $39\%$ $0.5\%$ $2\%$ $4\%$ $11\%$
PEOA $13\%$ $9\%$ $0.8\%$ $0.5\%$ $\% 33\%$
[…]
@@
[…]
@@
Method Max-FES $\Omega_{m0}] PREC$ $H_{0}$] PREC[$\sigma_{8}$] Time (s)
@@
It is worth noting that the results of each method are consistent in all cases, though the constraints are not centered relative to the expected values [^5] This can be addressed by including the covariance in the analysis, since our mock data sets are based on correlated observations that were treated independently.. Regardless, the overall offset is consistent for all of the methods and for this work important observation is the performance of EA with respect to MCMC. The whisker plot supports the sentiments on accuracy and precision of EA that have been reiterated throughout the results and the discussion. Note that when the full covariances of the data sets are considered, we can expect that generally the posteriors will be broader and in the case of very stringent data sets potential systematic biases on parameter estimation can be mitigated.
@@ 7 Conclusions
In this work, we introduced two novel evolutionary algorithms (EAs) for cosmological parameter estimation: the Improved Multi-Operator Differential Evolution (IMODE) and the Philippine Eagle Optimization Algorithm (PEOA). Among these, PEOA demonstrated performance comparable to the standard Markov Chain Monte Carlo (MCMC) method while surpassing other EAs, such as IMODE and the genetic algorithm (GA). These results underscore the potential of PEOA in cosmological applications, although it is crucial to interpret such findings within the specific context of parameter estimation. In optimization problems, there is no universally superior method; thus, this study serves as an initial benchmark for exploring EAs in cosmology with full systematics.
@@
The motivation for exploring alternative statistical methods is not to replace existing techniques but to complement them. This is particularly relevant in cosmology, where diverse data sets—ranging from supernovae and the cosmic microwave background to baryon acoustic oscillations and gravitational waves—must be reconciled under a consistent physical framework. Similarly, the application of diverse statistical methods, including EAs, can provide robust cross-validation and novel perspectives. Granted, the present work has no cosmological consequences and dealt with mock data. Nonetheless, it represents a step toward establishing a more diverse set of tools for confronting cosmological data.
Future research directions naturally emerge from this study. First, while GA has been applied to real cosmological data, the potential of IMODE and PEOA, among a full universe of nature-inspired EAs, remains to be tested on such data sets, which include full covariance structures. This will offer insights into their performance in more realistic scenarios. Additionally, EAs could play a significant role in cosmological model selection and in reconstructing cosmological functions such as spatial curvature and the dark energy density. Another promising avenue lies in leveraging the ability of EAs to explore high-dimensional and complex parameter spaces to investigate potential systematic effects in the data. Through simplified mock data sets, this work has opened up a path to a field of biology-inspired optimization techniques (beside the few well knowns such as genetic algorithm and particle swarm optimization) for cosmology, but the simplicity has also concealed EAs full potential. Fully tapping into EAs mathematical capabilities such as exploring multi-modal functions or likelihoods remains a worthy consideration in a future work, in cosmological parameter estimation and in other astrophysical contexts. On a practical side, we are preparing user-friendly python implementations of IMODE and PEOA to facilitate their adoption to cosmology and astrophysics.
Acknowledgements
@@ Appendix A Pseudo-algorithms
  • 2.
    $X_{\textrm{min}},X_{\textrm{max}}]$: search space
  • 3.
@@ A.1 Genetic Algorithm
  • 1.
    PSPS: population size
  • 2.
    crcr: crossover rate
  • 3.
@@
0:$f$, $D$, PSPS, lower bound $\mathbf{X}_{\min}$, upper bound $\mathbf{X}_{\max}$, crcr, $m$
[…]
1: Generate initial random population of size PSPS.;
@@
17:  The new generation is the collection of the elites and the offspring while maintaining the population size PSPS.
@@ A.2 Improved Multi-Operator Differential Evolution Algorithm
1: Define $Prob_{ls}\leftarrow 0.1$, $Max\text{-}FES$, $prob_{1}\leftarrow 1$, $prob_{2}\leftarrow 1$, NPNP, $G\leftarrow 1$, and $FES\leftarrow 0$;
2: Generate an initial random population $(X)$ of size NPNP;
@@
7:  Generate new population using the assigned DE operators, i.e., each operator opop evolves its assigned number of individuals $NP_{op}$;
8:  Calculate the diversity obtained from each operator opop $\left(D_{op}\right)$ and the quality rate of solutions $QR_{op}$;